If θ is a second quadrant angle and sinθ=15/17 , then find an exact value for
2sinθ+ 3cosθ
Thanks

Answer by lwsshak3(11628)

If θ is a second quadrant angle and sinθ=15/17 , then find an exact value for
2sinθ+ 3cosθ
***
sinx=15/17
cosx=-√1-(15/17)^2)=-√274/17
2sinx+3cosx=30/17-3√274/17=3(10-√274)/17